The following information applies to the questions displayed below.] Sanyu Sony started a new business and completed these transactions during December. Dec. 1 Sanyu Sony transferred $68,100 cash from a personal savings account to a checking account in the name of Sony Electric. 2 The company rented office space and paid $1,000 cash for the December rent. 3 The company purchased $14,300 of electrical equipment by paying $6,100 cash and agreeing to pay the $8,200 balance in 30 days. 5 The company purchased office supplies by paying $700 cash. 6 The company completed electrical work and immediately collected $1,200 cash for these services. 8 The company purchased $2,530 of office equipment on credit. 15 The company completed electrical work on credit in the amount of $4,500 18 The company purchased $400 of office supplies on credit. 20 The company paid $2,530 cash for the office equipment purchased on December 8. 24 The company billed a client $900 for electrical work completed; the balance is due in 30 days. 28 The company received $4,500 cash for the work completed on December 15. 29 The company paid the assistants salary of $2,000 cash for this month. 30 The company paid $550 cash for this months utility bill. 31 Sanyu Sony withdrew $940 cash from the company for personal use.Equity Liabilities Accounts Payable S. Sony Withdrawals Assets S. Sony, Capital RevenuesExpenses Office + Supplies Office Equipment Electrical Equipment + Accounts Date Cash Receivable Bal Bal Bal Bal Bal Bal Bal Bal Bal Bal Bal Bal Bal
Answer #1
Date Assets Liabilities Equity
Dec. Cash +

Accounts

Receivable

+ Office Supplies + Office Equipment + Electronic Equipment = Accounts Payable + S. Sony Capital -

S. Sony,

withdrawal

+ Revenue - Expenses
01 68,100 = + 68,100 - 1,000
02 (1,000) =
Bal. 67,100 = 68,100 - 1,000
03 (6,100) + 14,300 = 8,200 + 68,100
Bal. 61,000 + 14,300 = 8,200 + 68,100 - 1,000
05 (700) + 700
Bal. 60,300 + 700 + 14,300 = 8,200 + 68,100 - 1,000
06 1,200 + 1,200
Bal. 61,500 + 700 + 14,300 = 8,200 + 68,100 + 1,200 - 1,000
08 + 2,530 = 2,530
Bal. 61,500 + 700 + 2,530 + 14,300 = 10,730 + 68,100 + 1,200 - 1,000
15 + 4,500 + 4,500
Bal. 61,500 + 4,500 + 700 + 2,530 + 14,300 = 10,730 + 68,100 + 5,700 - 1,000
18 + 400 = 400
Bal. 61,500 + 4,500 + 1,100 + 2,530 + 14,300 = 11,130 + 68,100 + 5,700 - 1,000
20 (2,530) = (2,530)
Bal. 58,970 + 4,500 + 1,100 + 2,530 + 14,300 = 8,600 + 68,100 + 5,700 - 1,000
24 + 900 + 900
Bal. 58,970 + 5,400 + 1,100 + 2,530 + 14,300 = 8,600 + 68,100 + 6,600 - 1,000
28 4,500 + (4,500) =
Bal. 63,470 + 900 + 1,100 + 2,530 + 14,300 = 8,600 + 68,100 + 6,600 - 1,000
29 (2,000) = - 2,000
Bal. 61,470 + 900 + 1,100 + 2,530 + 14,300 = 8,600 + 68,100 + 6,600 - 3,000
30 (550) = - 550
Bal. 60,920 + 900 + 1,100 + 2,530 + 14,300 = 8,600 + 68,100 + 6,600 - 3,550
31 (940) = - 940
Bal. 59,980 + 900 + 1,100 + 2,530 + 14,300 = 8,600 + 68,100 - 940 + 6,600 - 3,550

These are called Accounting Equation. In these equations both sides are always equal.